A rigid vessel contains water at 400 °F. Heat is to be added so that the water passes through the critical point. What should the quality be at the temperature of 400° F.
1
Expert's answer
2020-10-20T14:11:19-0400

Firstly we will change the temprature into degree Celsius

"T(^oC)=\frac{5}{9}(T^oF)-32=\frac{5}{9}(400)-32=204.4 ^o C"

At this temperature water will not present it will change into stem and we know that

critical point for the steam is "(T_c=374 ^oC)"

From steam table we can find the value of specific volume at both temperature which will be important in finding the quality of steam

At Temperature T1= 204.4 oC


vf=0.001164 "\frac{m^3}{kg},v_g=0.11508 \frac{m^3}{kg}" , and volume of steam at critical point = 0.003105 "\frac{m^3}{kg}"


Quality of steam

"v_1=v_f+x(v_g-v_f)"

0.003105=0.001164+ x(0.11508-0.001164)

x=0.0171 It is the quality of steam